# An Iconic Heavy Hitter Algorithm Made Private

### v1

@@ Appendix B Deferred Proofs Section 4
    The counts of the items in the intersection remain unchanged and we remain in $(\mathsf{S2})$. Second, if $C_{t-1}[x^{*}]\neq C_{t-1}^{\prime}[z^{\prime}]$, the condition on $C_{t-1}[z]$ implies that $C_{t-1}[z]=C_{t-1}[x^{*}]=C_{t-1}^{\prime}[z^{\prime}]+1$. Therefore, we observe both
    $C_{t}[x_{t}]=C_{t-1}[x^{*}]+1=C_{t-1}^{\prime}[z^{\prime}]+2=C_{t}^{\prime}[x_{t}]+1,$
    which is condition $(\mathsf{S3})\text{(b)}$, and
    $C_{t}[z]=C_{t-1}[z]=C_{t-1}[x^{*}]=C_{t}^{\prime}[x^{*}],$
    which is condition $(\mathsf{S3})\text{(c)})$. Therefore, the process transitions to $(\mathsf{S3})$
@@
Appendix C Deferred Proofs Section 5
@@
Define the “stable envelope” event $\xi$ as the event that no item arrival in either stream exceeds the error envelope. That is, for
$E\;:=\;\bigl[\,\gamma_{1}(t)\ \geq\ \tilde{f}_{t}(x_{t})-f_{t}(x_{t})\ \geq\ \gamma_{2}(t)\qquad\forall x_{t}\in X\,\bigr]$
$E^{\prime}\;:=\;\bigl[\,\gamma_{1}(t)\ \geq\ \tilde{f}_{t}(x_{t}^{\prime})-f_{t}(x_{t}^{\prime})\ \geq\ \gamma_{2}(t)\qquad\forall x_{t}^{\prime}\in X^{\prime}\,\bigr],$
$\xi=E\land E^{\prime}$. Using a union bound on each stream arrival and both streams, by assumption, $\Pr[\xi]\geq 1-\delta$.
As the threshold is set at $\tau>T/\tilde{k}+\gamma_{1}(T)+\gamma_{2}(T)$, by Lemma 13, condition on event $\xi$, the released index sets $\hat{\mathcal{T}}$ and $\hat{\mathcal{T}}^{\prime}$ are contained in the intersection $\mathcal{T}\cap\mathcal{T}^{\prime}$. Let $m:=|\mathcal{T}\cap\mathcal{T}^{\prime}|\geq\tilde{k}-2$, and fix an arbitrary but common ordering of the labels in $\mathcal{T}\cap\mathcal{T}^{\prime}$. Define the vector-valued function $F(X)\in\mathbb{R}^{m}$ as the ordered list of frequency estimates. $F(X):=(\,\mathcal{D}.\mathsf{query}(x)\,)_{x\in\mathcal{T}\cap\mathcal{T}^{\prime}},$ and similarly $F(X^{\prime}):=(\,\mathcal{D}^{\prime}.\mathsf{query}(x)\,)_{x\in\mathcal{T}\cap\mathcal{T}^{\prime}}$. By assumption, for all measurable $A\subseteq\mathbb{R}^{m}$

### v2

@@ C.1 Lemma 4 Components
    The counts of the items in the intersection remain unchanged and we remain in $(\mathsf{S2})$. Second, if $C_{t-1}[x^{*}]\neq C_{t-1}^{\prime}[z^{\prime}]$, the condition on $C_{t-1}[z]$ implies that $C_{t-1}[z]=C_{t-1}[x^{*}]=C_{t-1}^{\prime}[z^{\prime}]+1$. Therefore, we observe
    $C_{t}[x_{t}]=C_{t-1}[x^{*}]+1=C_{t-1}^{\prime}[z^{\prime}]+2=C_{t}^{\prime}[x_{t}]+1\quad(\text{condition}(\mathsf{S3})\text{(b)})$
    and
    $C_{t}[z]=C_{t-1}[z]=C_{t-1}[x^{*}]=C_{t}^{\prime}[x^{*}]\quad(\text{condition}(\mathsf{S3})\text{(c)}).$
    Therefore, the process transitions to $(\mathsf{S3})$
@@
C.2 Performance and Utility
See 2
Proof.
Set $\tilde{k}=2k$ and run $\mathsf{SpaceSaving}$ with capacity $\tilde{k}$. This uses $\mathcal{O}(\tilde{k})=\mathcal{O}(k)$ words and requires $\mathcal{O}(1)$ update time using a standard implementation [8]. Let $C$ be the terminal (non-private) counters and add independent noise $\eta_{x}\sim\mathtt{Laplace}(1/\varepsilon)$, for each item $x\in\mathcal{T}$. Release labels with $C[x]+\eta_{x}>\tau$. As
$\frac{|X|}{k}-\gamma=2\cdot\frac{|X|}{2k}-\gamma>\frac{|X|}{2k}+\gamma+1,$
where the inequality follows by assumption, it follows that $\tau=\tfrac{|X|}{k}-\gamma$. By the error guarantee of $\mathsf{SpaceSaving}$ (Lemma 3),
$f_{x}\;\leq\;C[x]\;\leq\;f_{x}+\frac{|X|}{\tilde{k}}=f_{x}+\frac{|X|}{2k},$
for all $x\in\mathcal{T}$. For $\eta_{x}\sim\mathrm{Laplace}(1/\varepsilon)$ we have $\Pr[|\eta_{x}|>t]=e^{-\varepsilon t}$. Taking $t=\tfrac{1}{\varepsilon}\ln\left(\tfrac{1}{\delta}\right)$ gives, with probability $\geq 1-\delta$,
which matches the stated bound.
If $f_{y}>|X|/k$, then $C[y]\geq f_{y}>|X|/k$, so, with probability $1-\delta$,
$C[y]-\tau\;>\;\frac{|X|}{k}-\Big(\frac{|X|}{k}-\gamma\Big)=\gamma\;>\;0$
by the premise $|X|/(2k)>\gamma$. Therefore, $y$ fails the threshold only if $\eta_{y}<-\gamma$, which occurs with probability $\leq\tfrac{1}{2}e^{-\varepsilon\gamma}<\delta$. ∎
Appendix D Deferred Proofs Section 5
@@
Define the “stable envelope” event $\xi$ as the event that no frequency query exceeds the error envelope during the execution of the algorithm. That is, for
$E_{1}\;$ $:=\;\bigl[\,\gamma_{1}(t)\ \geq\ \tilde{f}_{t}(x_{t})-f_{t}(x_{t})\ \geq\ \gamma_{2}(t),\forall x_{t}\in X\,\bigr]$
$E_{1}^{\prime}\;$ $:=\;\bigl[\,\gamma_{1}(t)\ \geq\ \tilde{f}_{t}(x_{t}^{\prime})-f_{t}(x_{t}^{\prime})\ \geq\ \gamma_{2}(t),\forall x_{t}^{\prime}\in X^{\prime}\,\bigr]$
$E_{2}\;$ $:=\;\bigl[\,\gamma_{1}(T)\ \geq\ \tilde{f}_{T}(x)-f_{T}(x)\ \geq\ \gamma_{2}(T),\forall x\in\mathcal{T}\,\bigr]$
$E_{2}^{\prime}\;$ $:=\;\bigl[\,\gamma_{1}(T)\ \geq\ \tilde{f}_{T}(x^{\prime})-f_{T}(x^{\prime})\ \geq\ \gamma_{2}(T),\forall x^{\prime}\in\mathcal{T}^{\prime}\,\bigr]$
$\xi=E_{1}\land E_{1}^{\prime}\land E_{2}\land E_{2}^{\prime}$. There are at most $T+\tilde{k}$ frequency queries during algorithm execution. Thus, using a union bound across both streams, by assumption, $\Pr[\xi]\geq 1-\delta$.
The threshold is set at $\tau>T/\tilde{k}+2\gamma_{1}(T)+\gamma_{2}(T)$. Each estimate during post-processing has error at most $\gamma_{1}(T)$. By Lemma 11, an item in the symmetric difference has error at most
$\hat{f}_{T}(x)\leq f_{T}(x)+\gamma_{1}(T)\leq\frac{t}{\tilde{k}}+\gamma_{2}(T)+2\gamma_{1}(T)\leq\tau,$
where $t\leq T$ is the last time it was removed from the tracked set in the opposing stream. Therefore, conditioned on event $\xi$, the released index sets $\hat{\mathcal{T}}$ and $\hat{\mathcal{T}}^{\prime}$ are contained in the intersection $\mathcal{T}\cap\mathcal{T}^{\prime}$. Let $m:=|\mathcal{T}\cap\mathcal{T}^{\prime}|\geq\tilde{k}-2$, and fix an arbitrary but common ordering of the labels in $\mathcal{T}\cap\mathcal{T}^{\prime}$. Define the vector-valued function $F(X)\in\mathbb{R}^{m}$ as the ordered list of frequency estimates. $F(X):=(\,\mathcal{D}.\mathsf{query}(x)\,)_{x\in\mathcal{T}\cap\mathcal{T}^{\prime}},$ and similarly $F(X^{\prime}):=(\,\mathcal{D}^{\prime}.\mathsf{query}(x)\,)_{x\in\mathcal{T}\cap\mathcal{T}^{\prime}}$. By assumption, for all measurable $A\subseteq\mathbb{R}^{m}$